Mike paints a room in 5 hours. Mary paints the same room in 4 hours. Working together how long will it take them?

Answer by richwmiller(17219)

Let Mary do it all. Mike has more important things to do.
x/5+x/4=1
4x+5x=20
9x=20
x=20/9 2 2/9 hours.
They only gain 1 7/9 hours.
It would be a waste of time for Mike to travel over there to help her. He could get another job done instead of hauling equipment around.